Manuel Kauers Hagenberg, May 4, 2003 RISC-Linz

Tenses I

The Past Perfect

1 Participles

• Def. Let Xen be a regular German verb. Then Xend is called Partizip Präsens Aktiv (PPA) and geXt is called Partizip Perfekt Passiv (PPP) of Xen.

• Examples:

Infinitiv PPA  PPP
sagen  sagend  gesagt
hören  hörend  gehört
lachen  lachend  gelacht
lieben  liebend  geliebt
feiern  feiernd  gefeiert
arbeiten arbeitend gearbeitet

• PPA and PPP roughly correspond to the English -ing and -ed forms.

• Partizip forms can be used as adjectives, they are declinated accordingly.

• Examples:

Das gesagte Wort Das lachende Kind

Das gesagte Wort des lachenden Kindes

• Problem: Many verbs are not regular. Many irregular PPP forms end in -en instead of -t or the ge- is omitted. Many also change the root (Ablaut), and some are entirely out of rule.

• Examples:

Infinitiv PPP
kommen gekommen
sehen  gesehen
lesen  gelesen
gehen  gegangen
schreiben geschrieben
nehmen genommen
sein  gewesen
beginnen begonnen
vergessen vergessen

• In a splitting verb like an|kommen (arrive), the PPP’s ge- is put between the prefix and the root: angekommen.

Exercise

• What is the PPP to teilnehmen (participate)?

• Guess the infinitives of the following irregular PPPs

gegessen, getrunken, gegangen, angefangen, gewußt

• In the table of the irregular PPPs above, the PPA forms are missing which are regular. Add them. (Note that there is no PPA to sein.)

Infinitiv PPA

kommen sehen gehen schreiben nehmen beginnen vergessen

Past Perfect

• One way to express that something happened in the past is to use Perfekt.

• Perfekt is composed of a form of sein or haben as the conjugated part, and of the PPP form which goes to the end.

• Examples

das Buch – book

Gestern habe ich ein Buch über kommutative Algebra gelesen. Ich bin gestern mit dem Bus nach Linz gefahren.

• In a subclause, the conjugated part of the verb goes to the end, as usual. The participle stands in front of it.

• Example:

nachdem – after zurück – back

Nachdem ich am SFB-Seminar teilgenommen habe, fahre ich nach Hagenberg zurück.

• Whether a perfect is build with sein or haben depends on the verb. Here is a table:

with sein with haben
sein  haben
kommen nehmen
gehen  wissen
…  lesen

…

• As a rule of thumb, sein is used with verbs that don’t carry objects.

Exercise

• Transform the following sentences from present tense to past tense

Bruno schreibt Bücher über Gröbnerbasen. Ich vergesse das nicht. Kommst Du nach Strobl mit? Sag mir, wann Du nach Hause gehst!^{1}

• Make a collection of verb which you know and guess if their Perfekt is build with sein or haben.

1Only convert the subclause.